a solid 6.45 - cm square cube is made with a material that has a density of 1.13 g•cm^-3 . Calculate the centigram mass of this cube. Recall that the volume of a cubic shape is defined as the product of length x width x height


1
Expert's answer
2022-02-06T11:20:26-0500

Solution:

ρ = 1.13 g cm-3

s = 6.45 cm


Calculate the volume of cube:

V = s3 = (6.45 cm)3 = 268.336 cm3


Calculate the mass of cube:

m = ρ × V

m = (1.13 g cm-3) × (268.336 cm3) = 303.22 g


Convert grams (g) to centigrams (cg):

1 g = 100 cg

Hence,

m = (303.22 g) × (100 cg / 1 g) = 30322 cg

m = 30322 cg


Answer: 30322 cg